OpenStudy (anonymous):

factor 21x^2 + 43x - 14

OpenStudy (anonymous):

(3x+7)(7x-2) :)

OpenStudy (anonymous):

what would it be if set to zero?

OpenStudy (anonymous):

-7/3 and 2/7 \[3x+7=0, 3x=-7, x=-7/3, and 7x-2=0,7x=2,x=2/7\]